OpenStudy (anonymous):

the speed of a stream is 5mph. if a boat travels downstream in the same time that it takes to travel 36 miles upstream. what is the speed of the boat in still water?

OpenStudy (anonymous):

205/31

OpenStudy (anonymous):

typo* if a boat travels 72 miles downstream in the same time that it takes to travel 36 miles upstream

OpenStudy (anonymous):

15 mph

OpenStudy (anonymous):

t=d/r Solve the following for b, the boat speed in still water:\[\frac{36}{b-5}=\frac{72}{b+5}\]\[b=15 \text{ mph} \]